### v1

Abstract
The best known fully retroactive priority queue costs $O(\log^{2}m\allowbreak\log\log m)$ time per operation, where $m$ is the number of operations performed on the data structure. In contrast, standard (non-retroactive) and partially retroactive priority queues cost $O(\log m)$ time per operation. So far, it is unknown whether this $O(\log m)$ bound can be achieved for fully retroactive priority queues.
In this work, we study a restricted variant of priority queues known as monotonic priority queues. We show that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem. We design a fully retroactive monotonic priority queue with a cost of $O(\log m+T(m))$ time per operation, where $T(m)$ is the maximum between the query and the update time of a specific range-searching data structure with $m$ elements. Finally, we design a fully retroactive monotonic priority queue that costs $O(\log m\log\log m)$ time per operation.
keywords:
computational geometry, design and analysis, monotone priority queue, range tree, retroactive data structure, temporal data
category:
[…]
1 Introduction
Standard data structures operate only in the ?present?. That is, they only take into consideration their current state and ?forget? about their past states. For the vast majority of problems, standard data structures are enough and are the best option. However, standard data structures are not as versatile as they can be:
  • Their past states cannot be queried.
  • Once an element is deleted, all the information about the element is lost.
  • If an operation is mistakenly performed in the past, it is usually not possible to efficiently undo or modify this operation in the present.
Retroactive data structures are a type of data structure that can handle these problems. They were introduced by Demaine et al. [6] and are capable of modifying its past states and propagating the consequences of this modification until its present state. Furthermore, they have been used to solve some problems such as cloning Voronoi diagrams [8] and finding the shortest path in a dynamic graph [11].
A data structure is called partially retroactive if it can modify the past, but can only make queries in the present. A data structure is called fully retroactive if it can modify and query the past. It is possible to use a bank system as an analogy to better understand the different types of retroactivity:
  • A ‘standard’ bank system (data structure) can make new transactions (updates) in the present and show (query) the current balance.
  • A ‘retroactive’ bank system can modify past transactions (e.g., correct an old deposit), with all subsequent balances automatically updated to reflect this historical change.
  • A ‘partially retroactive’ bank system can make these historical changes, but still shows only the current balance.
  • Finally, a ‘fully retroactive’ bank system can both modify past transactions and show the account balance at any time (that is, after any transaction).
Priority queues
Let $m$ be the number of operations performed in the data structure, Demaine et al. [6] designed a partially retroactive priority queue with $O(\log m)$ time per operation, matching the bounds of a binary heap [12]. In contrast, the most efficient fully retroactive priority queue so far has $O(\log^{2}m\log\log m)$ time per operation [7]. It is known that some fully retroactive data structures need to have a nonconstant multiplicative slowdown over its partially retroactive version [2], conditioned on well-believed conjectures. However, it is unknown whether this is the case for retroactive priority queues.
A monotonic priority queue is a restricted variant of priority queues, where the extracted elements are restricted to form a non-decreasing function over time. One of its applications is in Dijkstra’s algorithm, since the discovered vertices naturally follow this restriction [5]. Given that a monotonic priority queue assumes a more restricted behavior, it is usually easier to design it efficiently than a general priority queue. In this work, we study how to design an efficient fully retroactive monotonic priority queue.
Our contributions
  • We show that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem (Section˜3).
  • We present a fully retroactive monotonic priority queue with $O(\log m+T(m))$ time per operation, where $T(m)$ is the maximum between the query and the update time of a specific range-searching data structure with $m$ elements (Section˜4).
  • We present a fully retroactive monotonic priority queue with $O(\log m\allowbreak\log\log m)$ time per operation (Section˜5).
2 Preliminaries
In this section, we introduce the background necessary to understand the results presented in the next sections. In Subsection˜2.1, we present the definitions used during this work. In Subsection˜2.2, we present some important characteristics of retroactive priority queues in order to facilitate the understanding of the topic. Finally, in Subsection˜2.3, we list the assumptions and notation used in the rest of this work.
2.1 Definitions
Definition 2.1.
A priority queue can be defined as an abstract data type that maintains a collection of elements and supports the following operations:
  • $\operatorname{insert}(x)$: Inserts an element $x$.
  • $\operatorname{get-min}()$: Returns the minimum element.
  • $\operatorname{extract-min}()$: Extracts the minimum element.
Definition 2.2.
A monotonic priority queue is a priority queue with the following (equivalent) constraints:
  • The extracted elements form a nondecreasing sequence [3]. That is, after an element is extracted, only elements greater than or equal to it can be extracted.
  • Elements smaller than the last extracted cannot be inserted. [5].
Definition 2.3.
By adding the concept of retroactivity to a priority queue, we can define a retroactive priority queue as an abstract data type with the following operations [6]:
  • $\operatorname{Insert}(\operatorname{insert}(x),t)$: Inserts a $\operatorname{insert}(x)$ operation at time $t$.
  • $\operatorname{Delete}(\operatorname{insert}(x),t)$: Deletes a $\operatorname{insert}(x)$ operation at time $t$.
  • $\operatorname{Insert}(\operatorname{extract-min},t)$: Inserts a $\operatorname{extract-min}()$ operation at time $t$.
  • $\operatorname{Delete}(\operatorname{extract-min},t)$: Deletes a $\operatorname{extract-min}()$ operation at time $t$.
  • $\operatorname{GetMin}(t)$: Returns the minimum element that exists in the priority queue at time $t$.
Note that a partially retroactive priority queue (PR-PQ) has the constraint that the $\operatorname{GetMin}$ operation can be performed only at the present time. That is, $\operatorname{GetMin}(t)$ can be executed only if $t=\infty$. A fully retroactive priority queue (FR-PQ) does not have this constraint.
Definition 2.4.
We define fully retroactive monotonic priority queues (FR-MonoPQ) in the same way as FR-PQ, but with the monotonic constraints applied.
2.2 Characteristics of Retroactive Priority Queues
Ray Shooting Analogy
There is an analogy between retroactive priority queues and the vertical ray shooting problem [6] that can help to understand the behavior of a retroactive priority queue. More specifically, it is possible to represent a retroactive priority queue in a plane in the following way (see Figure˜1 for an illustration of this representation):
  • Let $k$ be the value of an element, $t_{ins}$ be its insertion time, and $t_{del}$ be the time it is extracted. Each inserted element is a segment that goes from point $(t_{ins},k)$ to point $(t_{del},k)$. If the element is never extracted from the priority queue, $t_{del}=\infty$.
  • Let $t_{em}$ be the time that an $\operatorname{extract-min}()$ operation occurs. An $\operatorname{extract-min}()$ operation is a vertical ray that starts at $(t_{em},-\infty)$ and extends upwards. This ray stops at the first line segment that it intersects. (Note that this segment is the segment of the smallest element that exists at time $t$. Therefore, this ray effectively finds the elements that the $\operatorname{extract-min}()$ operation extracts.)
  • A $GetMin(t)$ operation can be seen as returning the first segment that a vertical ray starting from $(t,-\infty)$ intersects, without modifying the segment as an $\operatorname{extract-min}()$ operation does.
(a) A retroactive priority queue with only $insert()$ operations.
(b) An $\operatorname{extract-min}()$ operation shortens the first intersected segment, forming ?L? shapes.
(c) Example with more $\operatorname{extract-min}()$ operations.
(d) $GetMin(t)$ operation returning $x$.
Figure 1: Geometric view of retroactive priority queues. Blue segments represent elements, red segments represent $\operatorname{extract-min}()$ operations, and green dotted segments represent $GetMin(t)$ queries.
Chain of Reaction
Using this representation, it becomes easy to see why a retroactive priority queue can be harder to implement efficiently than a standard one. Since once a retroactive update is made in the past, the extraction time of $O(m)$ elements can potentially change non-trivially. In order to propagate these changes, complex approaches are necessary, which can result in bad efficiency for the operations overall. Figure 2 shows an example of this. In this example, after the execution of the operation, a subset of elements had its extraction time changed. Therefore, all queries made after this insertion can potentially return a different value than what it would return before.
(a) Current representation.
(b) Representation after an $Insert(insert(x),t)$ operation.
Figure 2: Example of the chain of reaction in retroactive priority queues. Lines in green represent the that occur because of the operation. Dotted lines represent the state of the segment before the operation. Filled lines represent the state of the segment after the operation.
Note that, in the partially retroactive version, since the past cannot be queried, we do not need to maintain the state of the priority queue at all times. Thus, there is more freedom in how we can propagate the changes caused by the operation. We only need to guarantee that the state of the priority queue is correct in the present. In the fully retroactive version, the ability to query at any point in time makes the process of updating the data structure while maintaining efficient queries more challenging, since it is apparently needed to maintain a correct priority queue state at all times.
Retroactive Monotonic Priority Queues
We can also use the ray shooting analogy to represent a retroactive monotonic priority queue. We only need to add the monotonic restriction to our previous representation. To do that, we use the second constraint of Definition˜2.4: Elements smaller than the last extracted cannot be inserted.
Let $t_{em}$ be the time that an extraction operation occurs. Let $k$ be the element that was extracted by it. To represent the monotonic restriction, we add a rectangle $(t_{em},\infty)\times(-\infty,k)$ for each extraction operation. Figure˜3 shows an illustration of this representation.
In a retroactive monotonic priority queue, it is not possible to insert any new element if its segment starts inside any of these rectangles. This is true because the segment of any element smaller than an already extracted element would start inside of them.
Note that this representation is for building intuition only. We will use a different one (in Subsection˜3.3) in order to implement retroactive operations efficiently.
(a) Example with one extraction.
(b) Example with more extractions.
Figure 3: Illustration of monotonic retroactive priority queues. The gray rectangles represent areas where elements cannot be inserted in, because of the monotonic restriction.
2.3 Notation and Assumptions
The notation used throughout the text is as follows:
  • $E$ is the set of elements inserted into the priority queue.
  • $m$ is the total number of operations performed on the priority queue so far.
  • We call $val[k]$ the $k$th smallest element inserted into the priority queue.
  • We call $em[k]$ the $k$th earliest $\operatorname{extract-min}()$ operation.
  • $\operatorname{lastExtracted}(t)$ is the last extracted element before or at time $t$. When there is no extracted value in this interval, $\operatorname{lastExtracted}(t)=$-$\infty$.
  • $\operatorname{insertionTime}(x)$ is the time of insertion of an element $x$ into the priority queue.
  • $\operatorname{extractionTime}(x)$ is the time of extraction of an element $x$ from the priority queue.
We will list the assumptions that we make about the priority queues that we will be working with, in order to make them explicit and easier to reference during the proofs.
Assumption 2.5 (Domain).
For notational convenience, all elements and times are assumed to be real numbers. (However, the presented algorithms work for any totally ordered set that supports constant-time comparison, without affecting the asymptotic complexities.)
Assumption 2.6 (Unique Elements).
For simplicity, we assume that all elements inserted in the priority queue are distinct. (This assumption can be ignored if we do the following: Let $x$ be the element and $t$ be its insertion time. Instead of referring to $x$ directly, refer to it as $(x,t)$ instead. In this way, all elements become effectively unique again, since $t$ is unique (Assumption˜2.7).)
{observation}
Since all elements are unique, we have that $val[i]<val[j]$ for all $i<j$.
Assumption 2.7 (Unique Operation Times).
We assume that each operation on the priority queue occurs at a unique time.
{observation}
Since all operations occur at different times, we have that $em[i]$ is executed before $em[j]$ for all $i<j$.
Assumption 2.8 (Consistency).
We assume that only valid operations are performed. Specifically, no $\operatorname{extract-min}()$ operation is performed at a time when the priority queue is empty, and no operation breaks the monotonic property of the priority queue after its execution.
Assumption 2.9 (Existence Boundaries).
We assume that if an element is extracted at time $t$, then the element does not exist at time $t$. In contrast, if an element is inserted at time $t$, then it is present in the priority queue at time $t$.
3 Key Ideas
In this section, we will prove things about monotonic priority queues that will be helpful in order to design a FR-MonoPQ (fully retroactive monotonic priority queue) in the next section. More specifically, we find a way to efficiently check if an element exists at any time (Subsections˜3.1 and 3.2) and use that to reduce the query of a retroactive monotonic priority queue to the range-searching problem (Subsection˜3.3).
3.1 Conditions of Existence
In order to implement the $\operatorname{GetMin}(t)$ operation, we need to return the minimum element that is present in the priority queue at time $t$. Therefore, $\operatorname{GetMin}(t)$ returns $\min\left\{{x\mid x\in E,\,x\text{ exists at time }t}\right\}$. Since finding the minimum depends on first checking if an element exists, we will focus on that later in Subsection˜3.3. For now, we will focus on efficiently checking if an element exists.
To check if an element exists at time $t$, we can use the following observations: {observation} After an element is inserted into the priority queue, the only way it can stop existing is by being extracted.
{observation}
An element $x$ exists only from its $\operatorname{insertionTime}(x)$ until just before its $\operatorname{extractionTime}(x)$. That is, $x$ only exists at times $t\in[\operatorname{insertionTime}(x),\operatorname{extractionTime}(x))$.
{remark*}
Subsection˜3.1 is true because the only operation that modify the priority queue other than insertion is the $\operatorname{extract-min}()$ operation. Furthermore, since all inserted elements are distinct; the only way for them to stop existing is to be extracted; and they exist at time of insertion but not at time of extraction (Assumption˜2.9), it follows that Subsection˜3.1 is also true.
Using Subsection˜3.1, it is possible to check if an element exists at time $t$. So, $\operatorname{GetMin(t)}$ returns
$\min\left\{{x\mid x\in E,\,\operatorname{insertionTime}(x)\leq t<\operatorname{extractionTime}(x)}\right\}.$
But, this is not very useful if we want to support retroactive operations, since one retroactive operation could change the extraction time of $O(m)$ elements, and we would need to update all of them (see Subsection˜2.2).
Therefore, we need a way to check if an element exists that depends on values that are easy to maintain when a retroactive operation occurs. Because of that, we show the following:
Lemma 3.1 (Conditions for Existence).
In a monotonic priority queue, an element $x$ exists at time $t$ iff $\operatorname{insertionTime}(x)<=t$ and $x>\operatorname{lastExtracted}(t)$.
Proof 3.2.
We first prove the forward direction, and then the backward direction. For both directions, assume it is a monotonic priority queue.
Forward Direction (Necessary Condition):
If $x$ exists at time $t$, then $\operatorname{insertionTime}(x)<=t$. This is true because if $x$ was inserted after time $t$, it would not exist at time $t$ yet (Subsection˜3.1).
Assume $x$ exists at time $t$. Assume for a contradiction that $x\leq\operatorname{lastExtracted}(t)$. There are two cases (== and <<). We show that a contradiction occurs in both.
Case 1: $x=\operatorname{lastExtracted}(t)$. Since $\operatorname{lastExtracted}(t)$ was extracted, it does not exist at time $t$. And $x$ exists at time $t$. Thus, they have the same value but are not the same element, contradicting that all elements are distinct (Assumption˜2.7).
Case 2: $x<\operatorname{lastExtracted}(t)$. Since, at time $t$, $\operatorname{lastExtracted}(t)$ was extracted and $x$ was not, $x$ will be extracted after a greater element was extracted, contradicting that it is a monotonic priority queue.
Therefore, if $x$ exists at time $t$, then $x>\operatorname{lastExtracted}(t)$. This completes the forward direction.
Backward Direction (Sufficient Condition):
Assume that $x>\operatorname{lastExtracted}(t)$. Assume that $\operatorname{insertionTime}(x)<=t$.
Assume for a contradiction that $x$ does not exist at time $t$. By Subsection˜3.1, the only way for it to not exist at time $t$ is if $\operatorname{extractionTime(x)}\leq t$ is true. However, this contradicts the assumption that it is a monotonic priority queue: Since $\operatorname{lastExtracted}(t), xx$ is not $\operatorname{lastExtracted}(x)$ and is greater than it. This implies that $x$ was extracted at some moment and then $\operatorname{lastExtracted}(x)$ was extracted after, creating a decreasing function over time. This completes the backward direction, and hence the proof.
Directly from Lemma˜3.1, we have a new way of finding the minimum element in the priority queue at any time.
Corollary 3.3 (GetMin Expression).
The $\operatorname{GetMin}(t)$ operation returns
$\min\,\{\,x\mid x\in E,\,\operatorname{insertionTime}(x)\leq t,\,x>\operatorname{lastExtracted}(t)\,\}.$ (1)
3.2 Finding the Last Extracted
Now we show a way to efficiently find $\operatorname{lastExtracted}(t)$. For that, we can maintain the list of $\operatorname{extract-min}()$ operations in a binary search tree and find the last extraction until time $t$, via a predecessor search on $t$. However, we have no way yet of finding which element this operation extracts. Thus, we prove the following lemma:
Lemma 3.4.
In a monotonic priority queue, $val[k]$ can only be extracted by $em[k]$.
Proof 3.5.
This lemma will be proven in two parts. For both parts, assume it is a monotonic priority queue.
Part 1: $val[k]$ can be extracted by $em[k]$.
Assume $val[k]$ is extracted by $em[k]$. By Subsection˜2.3, $val[k]<val[l]$ for all $k<l$. By Subsection˜2.3, $em[k]$ is executed before $em[l]$ for all $k<l$.
If $val[k]$ is extracted by $em[k]$ for all possible $k$, then $val[k]$ will be extracted before $val[l]$ for all $k<l$. Therefore, the extracted values create a increasing function as $k$ increases (that is, as time passes). This satisfies the monotonic priority queue restriction, showing that it is possible for this to happen.
Part 2: if $j\neq k$, $val[k]$ cannot be extracted by $em[j]$.
For $j$ to be different from $k$, $j$ needs to be greater or smaller than $k$. Therefore, we prove this part by these two cases.
Case 1: $j<k$. Assume for a contradiction that $val[k]$ is extracted by $em[j]$ and $j<k$. Via the pigeonhole principle, there is an element smaller than $val[k]$ that needs to be extracted by $em[l]$ and $l\geq k$. Since this element is smaller than $val[k]$ and is being extracted after $val[k]$, the extracted values forms decreasing sequence at some point, which contradicts the assumption that it is a monotonic priority queue.
Case 2: $j>k$. Assume for a contradiction that $val[k]$ is extracted by $em[j]$ and $j>k$. Via the pigeonhole principle, there is an element greater than $val[k]$ that needs to be extracted by $em[l]$ and $l\leq k$. Since this element is greater than $val[k]$ and is being extracted before $val[k]$, the extracted values forms decreasing sequence at some point, which contradicts the assumption that it is a monotonic priority queue.
Since the $k$th element cannot be extracted by any $j$ smaller or greater than $k$, it cannot be different from $k$. Thus, this part of the proof is complete.
In Part 1, we proved that $val[k]$ can be extracted by $em[k]$. In Part 2, we proved that it cannot be extracted by any other $\operatorname{extract-min}()$ operation. Therefore, $val[k]$ can only be extracted by $em[k]$.
Since we can now find which element each $\operatorname{extract-min}()$ operation extracts, efficiently finding $\operatorname{lastExtracted}(t)$ becomes straightforward.
Corollary 3.6.
In a monotonic priority queue, there is a procedure that, given $t$, finds $\operatorname{lastExtracted}(t)$ using $O(\log m)$ time.
Proof 3.7.
Let all inserted elements be stored in an order-statistic tree $T_{el}$ [4, Sec. 17.1]. Let all extraction times be stored in an order-statistic tree $T_{em}$. To find $\operatorname{lastExtracted}(t)$, we can execute the following algorithm:
  • (1)
    Find the predecessor of $t$ in $T_{em}$, calling it $\mathit{lastOp}$.
  • (2)
    Find the order of $\mathit{lastOp}$ in $T_{em}$, calling it $k$.
  • (3)
    Find the $k$th smallest element of $T_{el}$, calling it $x$.
  • (4)
    Return $x$.
We claim that $x=\operatorname{lastExtracted}(t)$. The proof is as follows. In Step 1, we find $\mathit{lastOp}$, the last extraction operation until time $t$. (Note that $\mathit{lastOp}$ extracts $\operatorname{lastExtracted}(t)$. Hence, we need to return the element that $\mathit{lastOp}$ extracts.) In Step 2, we find its order $k$. Hence, $\mathit{lastOp}=em[k]$. We know that $\mathit{lastOp}=em[k]$ extracts $val[k]$ (Lemma˜3.4). Thus, in Step 3 we find $x=val[k]$. Since $x$ is the element extracted by $\mathit{lastOp}$, $x=\operatorname{lastExtracted}(t)$.
Finding the predecessor of a value, the order of a value, and the $k$th smallest element in an order-statistic tree can be done in $O(\log n)$ time each, where $n$ is the number of elements in it. Since the trees store all operations, they can have at most $m$ elements. Therefore, $n<=m$, and the algorithm takes $O(\log m)$ time in total.
{remark*}
We will not use it, but we can find $\operatorname{extractionTime}(x)$ in $O(\log m)$ time in a similar way: (1) Find the order of $x$, calling it $k$; (2) find the operation that extracts $x$ (by finding $em[k]$); (3) return the time that $em[k]$ is performed.
3.3 Reduction to Range Searching
Now that we know that $\operatorname{GetMin}(t)$ is equivalent to Expression 1 and that we can efficiently find $\operatorname{lastExtracted}(t)$, we just need to find the minimum between all elements that exist at time $t$.
To do that efficiently, we will represent the monotonic priority queue geometrically. The idea comes directly from Expression 1. We plot all the elements and their insertion times as points in a 2D plane. Once we do that, it can be noticed that the two conditions of existence of an element at time $t$ (Lemma˜3.1) become a rectangle intersection. The $\operatorname{GetMin}(t)$ query then becomes finding the lowest point that intersects with the rectangle. Figure˜4 shows an example of this representation. Here is a detailed specification:
Lemma 3.8.
We can represent a monotonic priority queue in the following way: All elements $e$ in $E$ become points $(\operatorname{insertionTime}(e),e)$ in a plane.
Let $R(t)$ be the 2-sided rectangle $(-\infty,t]\times(\operatorname{lastExtracted}(t),\infty)$. Let $p$ be the point with minimum $y$-coordinate that intersects with $R(t)$. In this representation, the $\operatorname{GetMin}(t)$ query returns the $y$-coordinate of $p$.
Proof 3.9.
For a point $(\operatorname{insertionTime}(e),\allowbreak e)$ to intersect with $R(t)$ it needs to have $\operatorname{insertionTime}(e)\leq t$ and $e>\operatorname{lastExtracted}(t)$, which are exactly the conditions of existence of an element at time $t$ (Lemma˜3.1). Therefore, all points that intersect $R(t)$ represent elements that exist at time $t$. And since the $y$-coordinate of a point stores the element it represents, the point with the minimum $y$-coordinate that is inside the rectangle represents the smallest element that exists at time $t$.
Figure 4: Illustration of Lemma˜3.8. The point that intersects the rectangle is $\operatorname{lastExtracted}(t)$. All points inside the shaded area (the rectangle) exist at time $t$, except $\operatorname{lastExtracted}(t)$. The marked point is the point with minimum value among all the points that exist at time $t$.
It turns out that the problem of getting information about the set of points within a given rectangle is a problem called range searching. This is a well-studied problem with many variations with its own specialized data structures for each of them (see [1] for a survey on the subject).
For our variation, we need to be able to insert and delete points (retroactive insertion). We also need to find the point with minimum $y$-coordinate that intersects with a given 2-sided orthogonal rectangle ($\operatorname{GetMin}(t)$ query). Therefore, we need the following data structure:
Definition 3.10.
Let $n$ be the number of points in the data structure, we define a min-y range-searching data structure as a data structure that maintains a collection of 2D points $P$ and has the following characteristics:
  • It can insert a new point into $P$, using $O(U(n))$ time.
  • It can remove an existing point from $P$, using $O(U(n))$ time.
  • Given a 2-sided orthogonal rectangle $R$, it can find the point with minimum $y$-coordinate inside $R$, using $Q(n)$ time.
  • It uses $S(n)$ space.
Because we have not found a data structure specialized for Definition˜3.10, for now we will not specify which data structure we will use to maintain generality. However, in Section˜5, we will extend an existing range-searching data structure so it satisfies Definition˜3.10.
4 The Data Structure
Using what we know from the previous section, we can now design an efficient FR-MonoPQ.
Theorem 4.1.
There is a FR-MonoPQ with the following costs:
  • $O(\log m)$ time per $\operatorname{Insert/Delete}(\operatorname{\operatorname{extract-min}()},t)$ operation.
  • $O(\log m+U(m))$ time per $\operatorname{Insert/Delete}(\operatorname{insert}(x),t)$ operation.
  • $O(\log m+Q(m))$ time per $\operatorname{GetMin}(t)$ operation.
  • $O(m+S(m))$ space.
Here $U(m)$, $Q(m)$, $S(m)$ are the update, query and space cost of a min-y range-searching data structure with $m$ elements, respectively.
Proof 4.2.
We use three auxiliary data structures $T_{el}$, $T_{em}$, $T_{ins}$:
  • $T_{el}$: An order-statistic tree that stores all inserted elements.
  • $T_{em}$: An order-statistic tree that stores all the times that an $\operatorname{extract-min}()$ operation occurs.
  • $T_{ins}$: A min-y range-searching data structure that stores all inserted elements and their insertion time. More specifically, let $(t,x)$ be an ordered pair that represents an element $x$ inserted at time $t$. $T_{ins}$ stores the pairs $(t,x)$ as points.
We first design the $\operatorname{GetMin}(t)$ operation. As we showed in Subsection˜3.3, the $\operatorname{GetMin}(t)$ operation can be solved by a query on a min-y range-searching data structure. Hence, we simply query $T_{ins}$. But to do that, we first have to find $\operatorname{lastExtracted}(t)$ to create the rectangle. Accordingly, this is how we can implement $\operatorname{GetMin}(t)$:
  • (1)
    Find $\operatorname{lastExtracted}(t)$.
  • (2)
    Query $T_{ins}$ with the rectangle $(-\infty,t]\times(\operatorname{lastExtracted}(t),\infty)$. Let $p$ be the returned point.
  • (3)
    Return the $y$-coordinate of $p$.
Finding $\operatorname{lastExtracted}(t)$ takes $O(\log m)$ time (Corollary˜3.6). And querying $T_{ins}$ takes $O(Q(m))$ time. Therefore, in total, $\operatorname{GetMin}(t)$ takes $O(\log m+Q(m))$ time.
Next, we design the other operations. Since $T_{el}$ and $T_{ins}$ store the elements, they need to be updated every time an $\operatorname{Insert}/\operatorname{Delete}(\operatorname{insert}(x),t)$ operation is performed. Accordingly, this is how we can implement $\operatorname{Insert}(\operatorname{insert}(x),t)$:
  • (1)
    Insert $x$ into $T_{el}$.
  • (2)
    Insert the pair $(t,x)$ into $T_{ins}$.
Note that Step 1 takes $O(\log m)$ time and Step 2 takes $O(U(m))$ time. Therefore, in total, this operation takes $O(\log m+U(m))$ time. The $\operatorname{Delete}(\operatorname{insert}(x),t)$ operation is implemented similarly, but it removes the element instead of inserting.
Since $T_{em}$ stores all extraction times, it needs to be updated every time an $\operatorname{Insert}/\allowbreak\operatorname{Delete}(\operatorname{extract-min},t)$ operation is performed. Accordingly, this is how we can implement $\operatorname{Insert}(\operatorname{extract-min},t)$: Insert $t$ into $T_{em}$.
Since this operation is just an insertion in $T_{em}$, it takes $O(\log m)$ time. And again, the $\operatorname{Delete}(\operatorname{extract-min},t)$ operation is implemented similarly, but it removes the time instead of inserting.
Finally, we do the space analysis. Note that each operation adds at most one element to each data structure. Hence, each data structure can have at most $m$ elements. Therefore, $T_{el}$ and $T_{ins}$ use $O(m)$ space each, $T_{ins}$ uses $S(m)$ space, and in total they use $O(m+S(m))$ space.
5 Choosing a Range-Searching Data Structure
In the previous sections, we presented a FR-MonoPQ that uses a generic min-y range-searching data structure. We did that because we have not found a range-searching data structure for our specific use case of Definition˜3.10. In this section, we show that the 2D range tree of Mehlhorn and Näher [10, Thm. 7] can be used to answer queries in the form: Given an orthogonal rectangle, find the point inside this rectangle that has the smallest $y$-coordinate.
The 2D range tree of Mehlhorn and Näher [10, Thm. 7] has the following characteristics:
  • It maintains a set of $n$ points $P$ in a plane.
  • It can add or remove a point in $O(\log n\log\log n)$ amortized time. (Dietz and Raman [9, Sec. 4.3] showed how to make this bound worst-case.)
  • All points are saved in a balanced primary tree, ordered by $x$-coordinate. Each node in the primary tree saves a collection of points ordered by $y$-coordinate. [^1] In their paper, they described a general technique and said that it can be applied in a straightforward way for range trees. So, they did not describe the tree structure explicitly, only its complexity. What we present here is our own interpretation of how the range tree can be organized. Each collection stores all the points inside the subtree of its node. [^2] Actually, each collection needs to store a superset of the points inside the subtree of its nodes. But, we stated otherwise for simplicity.
  • If a collection is found by walking on the primary tree from the root and $y$ is known before this walk starts, it is possible to find the successor of $y$ in the collection using $O(\log\log n)$ time [10, Lem. 3].
  • It uses $O(n\log n)$ space.
Here is the procedure for finding the point inside a given orthogonal rectangle with the smallest $y$-coordinate:
  • (1)
    Find all $O(\log n)$ collections of points that are within the horizontal range of the rectangle using the primary tree.
  • (2)
    For each of these collections of points, find the point with the smallest $y$-coordinate within the vertical range of the rectangle. (That is, find the successor of the smallest $y$-coordinate of the rectangle in each collection. If the found point is not inside the rectangle or does not exist, ignore it.)
  • (3)
    Of all the points found in Step 2, return the one with the smallest $y$-coordinate.
First, we check its correctness: Step 1 finds the set $S$ of points that are within the horizontal range of the rectangle. Then, Steps 2 and 3 find in $S$ the point with smallest $y$-coordinate that is in the vertical range of the rectangle. Therefore, this procedure finds the element with the smallest $y$-coordinate inside the rectangle.
Now we do the time analysis: Step 1 takes $O(\log n)$ time, since the primary tree is balanced. Step 2 takes $O(\log\log n)$ time [10, Lem. 3]. Since we have performed an operation that cost $O(\log\log n)$ time in $O(\log n)$ collections, in total the complexity of this procedure is $O(\log n\log\log n)$.
This data structure supports finding the point inside a given orthogonal rectangle with the smallest $y$-coordinate, and allows for insertion and deletion of points. Therefore, it satisfies the definition of a min-y range-searching data structure. Thus, using it directly in Theorem˜4.1, we obtain the following:
Corollary 5.1.
There is a FR-MonoPQ with the following costs:
  • $O(\log m)$ time per $\operatorname{Insert/Delete}(\operatorname{\operatorname{extract-min}()},t)$ operation.
  • $O(\log m\log\log m)$ time per $\operatorname{Insert/Delete}(\operatorname{insert}(x),t)$ operation.
  • $O(\log m\log\log m)$ time per $\operatorname{GetMin}(t)$ operation.
  • $O(m\log m)$ space.
Proof 5.2.
Theorem˜4.1 has the following bounds for a FR-MonoPQ:
  • $O(\log m)$ time per $\operatorname{Insert/Delete}(\operatorname{\operatorname{extract-min}()},t)$ operation.
  • $O(\log m+U(m))$ time per $\operatorname{Insert/Delete}(\operatorname{insert}(x),t)$ operation.
  • $O(\log m+Q(m))$ time per $\operatorname{GetMin}(t)$ operation.
  • $O(m+S(m))$ space.
Here $U(m)$, $Q(m)$, $S(m)$ are the update, query and space cost of a min-y range-searching data structure with $m$ elements, respectively.
We can use the data structure we described in this section directly in Theorem˜4.1, since it satisfies Definition˜3.10. Therefore, substituting $U(m)$ with $O(\log m\log\log m)$, $Q(m)$ with $O(\log m\log\log m)$, and $S(m)$ with $O(m\log m)$, we get the stated bounds.
6 Concluding Remarks
In this paper, we showed that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem. Then, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m+U(m))$ time per retroactive insertion, with $O(\log m+Q(m))$ time per retroactive query, and that uses $O(m\log m)$ space. Here $U(m)$, $Q(m)$ and $S(m)$ are the update, query, and space cost of a min-y range-searching data structure, respectively. Finally, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m\log\log m)$ time per retroactive query and insertion, and that uses $O(m\log m)$ space.
In future works, we intend to study the upper bound of other restricted versions of priority queues with full retroactivity, study the upper bound of the general fully retroactive priority queue, and/or study the lower bound of retroactive priority queues.

### v2

@@
 Abstract
-The best known fully retroactive priority queue costs $O(\log^{2}m\allowbreak\log\log m)$ time per operation, where $m$ is the number of operations performed on the data structure.
-In contrast, standard (non-retroactive) and partially retroactive priority queues cost $O(\log m)$ time per operation.
+The best-known fully retroactive priority queue costs $O(\log^{2}m\allowbreak\log\log m)$ time per operation, where $m$ is the number of operations performed on the data structure.
+In contrast, standard (non-retroactive) and partially retroactive priority queues can cost $O(\log m)$ time per operation.
 So far, it is unknown whether this $O(\log m)$ bound can be achieved for fully retroactive priority queues.
 In this work, we study a restricted variant of priority queues known as monotonic priority queues.
-We show that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem.
-We design a fully retroactive monotonic priority queue with a cost of $O(\log m+T(m))$ time per operation, where $T(m)$ is the maximum between the query and the update time of a specific range-searching data structure with $m$ elements.
+First, we show that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem.
+Then, we design a fully retroactive monotonic priority queue with a cost of $O(\log m+T(m))$ time per operation, where $T(m)$ is the maximum between the query and the update time of a specific range-searching data structure with $m$ elements.
 Finally, we design a fully retroactive monotonic priority queue that costs $O(\log m\log\log m)$ time per operation.
@@
 computational geometry, design and analysis, monotone priority queue, range tree, retroactive data structure, temporal data
+keywords:
+computational geometry, monotone priority queue, range tree, retroactive data structure
 category:
@@
 It is possible to use a bank system as an analogy to better understand the different types of retroactivity:
-  • A ‘standard’ bank system (data structure) can make new transactions (updates) in the present and show (query) the current balance.
-  • A ‘retroactive’ bank system can modify past transactions (e.g., correct an old deposit), with all subsequent balances automatically updated to reflect this historical change.
-  • A ‘partially retroactive’ bank system can make these historical changes, but still shows only the current balance.
-  • Finally, a ‘fully retroactive’ bank system can both modify past transactions and show the account balance at any time (that is, after any transaction).
+  • A standard bank system (data structure) can make new transactions (updates) in the present and show (query) the current balance.
+  • A retroactive bank system can modify past transactions (e.g., correct an old deposit), with all subsequent balances automatically updated to reflect this historical change.
+  • A partially retroactive bank system can make these historical changes but still shows only the current balance.
+  • Finally, a fully retroactive bank system can both modify past transactions and show the account balance at any time (that is, after any transaction).
 Priority queues
-Let $m$ be the number of operations performed in the data structure, Demaine et al.
+Let $m$ be the number of operations performed in the data structure.
+Demaine et al.
 [6] designed a partially retroactive priority queue with $O(\log m)$ time per operation, matching the bounds of a binary heap [12].
@@
   • The extracted elements form a nondecreasing sequence [3]. That is, after an element is extracted, only elements greater than or equal to it can be extracted.
-  • Elements smaller than the last extracted cannot be inserted. [5].
+  • Elements smaller than the last extracted cannot be inserted [5].
 Definition 2.3.
@@
 Definition 2.4.
-We define fully retroactive monotonic priority queues (FR-MonoPQ) in the same way as FR-PQ, but with the monotonic constraints applied.
+We define fully retroactive monotonic priority queues (FR-MonoPQ) in the same way as FR-PQ but with the monotonic constraints applied.
 2.2 Characteristics of Retroactive Priority Queues
@@
   • A $GetMin(t)$ operation can be seen as returning the first segment that a vertical ray starting from $(t,-\infty)$ intersects, without modifying the segment as an $\operatorname{extract-min}()$ operation does.
-(a) A retroactive priority queue with only $insert()$ operations.
+(a) A retroactive priority queue with only insert operations.
 (b) An $\operatorname{extract-min}()$ operation shortens the first intersected segment, forming ?L? shapes.
@@
 Since once a retroactive update is made in the past, the extraction time of $O(m)$ elements can potentially change non-trivially.
-In order to propagate these changes, complex approaches are necessary, which can result in bad efficiency for the operations overall.
+Thus, in order to propagate these changes, complex approaches are necessary, which can result in bad efficiency for the operations overall.
 Figure 2 shows an example of this.
@@
 Filled lines represent the state of the segment after the operation.
-Note that, in the partially retroactive version, since the past cannot be queried, we do not need to maintain the state of the priority queue at all times.
+Note that in the partially retroactive version, since the past cannot be queried, we do not need to maintain the state of the priority queue at all times.
 Thus, there is more freedom in how we can propagate the changes caused by the operation.
@@
 For notational convenience, all elements and times are assumed to be real numbers.
-(However, the presented algorithms work for any totally ordered set that supports constant-time comparison, without affecting the asymptotic complexities.)
-Assumption 2.6 (Unique Elements).
-For simplicity, we assume that all elements inserted in the priority queue are distinct.
-(This assumption can be ignored if we do the following: Let $x$ be the element and $t$ be its insertion time.
-Instead of referring to $x$ directly, refer to it as $(x,t)$ instead.
-In this way, all elements become effectively unique again, since $t$ is unique (Assumption˜2.7).)
-{observation}
-Since all elements are unique, we have that $val[i]<val[j]$ for all $i<j$.
+However, the presented algorithms work for any totally ordered set that supports constant-time comparison between two elements, without affecting the asymptotic complexities.
 Assumption 2.6 (Unique Operation Times).
@@
 Since all operations occur at different times, we have that $em[i]$ is executed before $em[j]$ for all $i<j$.
+Assumption 2.7 (Unique Elements).
+For simplicity, we assume that all elements inserted in the priority queue are distinct.
+This assumption can be ignored if we do the following: Let $x$ be the element and $t$ be its insertion time.
+Instead of referring to $x$ directly, refer to it as $(x,t)$ instead.
+In this way, all elements become effectively unique again, since $t$ is unique (Assumption˜2.6).
+{observation}
+Since all elements are unique, we have that $val[i]<val[j]$ for all $i<j$.
 Assumption 2.8 (Consistency).
@@
 Assumption 2.9 (Existence Boundaries).
-We assume that if an element is extracted at time $t$, then the element does not exist at time $t$.
-In contrast, if an element is inserted at time $t$, then it is present in the priority queue at time $t$.
+We assume that if an element is extracted at time $t$, then the element does not exist at time $t$. Also, if an element is inserted at time $t$, then it is present in the priority queue at time $t$.
 3 Key Ideas
@@
 An element $x$ exists only from its $\operatorname{insertionTime}(x)$ until just before its $\operatorname{extractionTime}(x)$.
-That is, $x$ only exists at times $t\in[\operatorname{insertionTime}(x),\operatorname{extractionTime}(x))$.
+That is, $x$ only exists at times $t\in[\operatorname{insertionTime}(x),\allowbreak\operatorname{extractionTime}(x))$.
 {remark*}
-Subsection˜3.1 is true because the only operation that modify the priority queue other than insertion is the $\operatorname{extract-min}()$ operation.
-Furthermore, since all inserted elements are distinct; the only way for them to stop existing is to be extracted; and they exist at time of insertion but not at time of extraction (Assumption˜2.9), it follows that Subsection˜3.1 is also true.
+Subsection˜3.1 is true because the only operation that modify the priority queue other than insertion is the $\operatorname{extract-min}()$ operation; and Subsection˜3.1 is true because the only way for the elements to stop existing is to be extracted (Subsection˜3.1), they exist at time of insertion but not at time of extraction (Assumption˜2.9), and all elements are distinct (Assumption˜2.7).
 Using Subsection˜3.1, it is possible to check if an element exists at time $t$. So, $\operatorname{GetMin(t)}$ returns
@@
 And $x$ exists at time $t$.
-Thus, they have the same value but are not the same element, contradicting that all elements are distinct (Assumption˜2.7).
+Thus, they have the same value but are not the same element, contradicting that all elements are distinct (Assumption˜2.6).
 Case 2: $x<\operatorname{lastExtracted}(t)$.
@@
 Since we can now find which element each $\operatorname{extract-min}()$ operation extracts, efficiently finding $\operatorname{lastExtracted}(t)$ becomes straightforward.
-Corollary 3.6.
+Lemma 3.6.
 In a monotonic priority queue, there is a procedure that, given $t$, finds $\operatorname{lastExtracted}(t)$ using $O(\log m)$ time.
@@
     Return the $y$-coordinate of $p$.
-Finding $\operatorname{lastExtracted}(t)$ takes $O(\log m)$ time (Corollary˜3.6).
-And querying $T_{ins}$ takes $O(Q(m))$ time.
-Therefore, in total, $\operatorname{GetMin}(t)$ takes $O(\log m+Q(m))$ time.
+Finding $\operatorname{lastExtracted}(t)$ takes $O(\log m)$ time (Lemma˜3.6). And querying $T_{ins}$ takes $O(Q(m))$ time. Therefore, in total, $\operatorname{GetMin}(t)$ takes $O(\log m+Q(m))$ time.
 Next, we design the other operations.
@@
 5 Choosing a Range-Searching Data Structure
-In the previous sections, we presented a FR-MonoPQ that uses a generic min-y range-searching data structure.
-We did that because we have not found a range-searching data structure for our specific use case of Definition˜3.10.
-In this section, we show that the 2D range tree of Mehlhorn and Näher [10, Thm.
+In this section, for the purpose of satisfying Definition˜3.10, we show that the 2D range tree of Mehlhorn and Näher [10, Thm.
 7] can be used to answer queries in the form: Given an orthogonal rectangle, find the point inside this rectangle that has the smallest $y$-coordinate.
+Later, we directly use this range tree in Theorem˜4.1 to obtain new bounds for the FR-MonoPQ.
 The 2D range tree of Mehlhorn and Näher [10, Thm. 7] has the following characteristics:
@@
 Thus, using it directly in Theorem˜4.1, we obtain the following:
-Corollary 5.1.
+Theorem 5.1.
 There is a FR-MonoPQ with the following costs:
@@
   • $O(m+S(m))$ space.
-Here $U(m)$, $Q(m)$, $S(m)$ are the update, query and space cost of a min-y range-searching data structure with $m$ elements, respectively.
 We can use the data structure we described in this section directly in Theorem˜4.1, since it satisfies Definition˜3.10.
 Therefore, substituting $U(m)$ with $O(\log m\log\log m)$, $Q(m)$ with $O(\log m\log\log m)$, and $S(m)$ with $O(m\log m)$, we get the stated bounds.
-6 Concluding Remarks
+6 Conclusion
 In this paper, we showed that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem.
-Then, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m+U(m))$ time per retroactive insertion, with $O(\log m+Q(m))$ time per retroactive query, and that uses $O(m\log m)$ space.
-Here $U(m)$, $Q(m)$ and $S(m)$ are the update, query, and space cost of a min-y range-searching data structure, respectively.
+Then, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m+U(m))$ time per retroactive insertion, with $O(\log m+Q(m))$ time per retroactive query, and that uses $O(m+S(m))$ space.
 Finally, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m\log\log m)$ time per retroactive query and insertion, and that uses $O(m\log m)$ space.